solution of the problem cubical fish tank having length,breadth and height 20,30 and 20 cm respectively...it is placed in a horizontal table and it is three quarters full of water..find the area of the tank which in contact with water
Answer by Theo(13342)

length = 20
breadth = 30
height = 20

tank is 3/4 filled.

length = 20
breadth = 30
height = 15

the length and breadth remains the same, but the height is now 15 instead of 20.

you are looking for the surface area of these new dimensions.

bottom = length * bread = 20 * 30 = 600 cm^2.

long side = height * breadth = 15 * 30 = 450 cm^2.

short side = height * length = 15 * 20 = 300 cm^2.

total surface area touching the water is equal to bottom + 2 * short side + 2 * long side = 600 + 900 + 600 = 2100 cm^2.